# The Heap: A Contamination-Free Multilingual Code Dataset for Evaluating Large Language Models

### v1

@@ Abstract
Index Terms:
Dataset, Evaluation, Large Language Models, Open Science, Data Contamination, Multilingual
@@ I Introduction
The data-intensive training process of Large Language Models (LLMs) has driven the release of numerous large-scale datasets, particularly for code, to facilitate the development of new models. This rapid increase in the amount of training data used to pre-train LLMs has resulted in extensive datasets covering almost all publicly available code [1, 2, 3].
To assess the success of such LLMs in downstream tasks, fresh data not seen during training is needed. Otherwise such evaluations are contaminated, possibly resulting in overly optimistic results. Unfortunately, obtaining such non-contaminated data is increasingly difficult. In fact, a recent study establishes that only 10% of investigations involving LLMs deduplicate their data with respect to the training data in order to avoid contamination [4].
To address this, we propose The Heap, a dataset of not previously used code that can be used for contamination-free multilingual evaluation of LLMs in downstream tasks. We address contamination in two ways. First, we select code with a non-permissive license, such as the GNU General Public License. Using such code for training is unattractive, as it may require the end user to publicly release all code in their code bases. Second, we pre-conduct computationally expensive near and exact deduplication, removing code that is used in other datasets widely used for training such as The Stack [1].
@@ II Collection
Using the search API, we collect our dataset from GitHub, a commonly used online platform for sharing code repositories. This collection process mimics the data distribution of other large-scale datasets [5, 3, 6, 7, 8], minimizing the probability of including confounding factors in the dataset, such as drifts in the representations of data [9].
@@ II-A Programming Languages
Ada 676 41,367 35,425
Agda 142 5,483 5,113
ANTLR 101 564 541
Apex 253 17,833 7,641
Assembly 7,100 208,896 104,901
C 50,000 16,585,280 4,960,192
C# 50,000 5,906,716 3,770,829
C++ 50,000 14,891,856 4,811,620
Clojure 27,107 380,567 273,181
Cobol 341 2,242 1,208
Common Lisp 796 45,083 16,968
Coq 477 54,137 26,175
Crystal 368 11,606 7,300
Cuda 1,191 26,948 13,359
D 1,185 185,630 126,111
Dart 11,907 484,935 413,203
EJS 1,475 15,513 12,884
Elixir 2,371 643,856 127,910
Emacs Lisp 377 8,260 7,963
Erlang 1,240 55,932 32,049
F# 876 22,152 16,015
Forth 222 28,287 7,932
Go 50,000 8,506,379 2,355,716
Groovy 2,198 60,299 48,353
Hack 1,379 84,916 37,405
Haskell 8,023 122,788 111,234
Java 50,000 6,989,601 5,197,338
JavaScript 50,000 8,289,901 3,393,747
Julia 2,859 46,284 38,381
Kotlin 21,665 1,467,343 1,045,396
Less 433 17,276 7,389
Lua 42,241 4,605,230 913,898
Mathematica 1,528 164,498 89,853
MATLAB 20,828 1,051,354 665,659
NetLogo 332 900 863
NewLisp 35 5,819 5,148
Nix 1,892 75,093 71,199
Objective-C 7,700 1,899,714 698,137
OCaml 1,961 121,890 69,171
Pascal 5,218 330,832 225,749
Perl 14,673 1,798,520 629,769
PHP 50,000 12,707,727 3,363,040
Processing 2,950 24,723 20,343
Prolog 1,071 38,995 20,279
Python 50,000 2,290,182 1,792,451
R 44,993 589,139 374,812
Raku 158 1,384 1,306
Ruby 13,378 1,579,655 794,364
Rust 42,847 2,496,177 844,258
Scala 5,893 749,370 224,021
Scheme 1,878 106,620 54,226
Scilab 199 4,531 4,084
SQL 130 47,185 41,178
Starlark 146 524 498
Swift 13,924 633,819 439,565
Vue 14,858 457,605 323,672
WebAssembly 68 834 587
Total 733,663 96,990,250 38,681,609
@@ II-B Query
We focus on repositories that have one of the targeted languages as the main language of the repository. We further select only repositories that are licensed under non-permissive licenses. We choose non-permissive licenses as an initial filter for repositories, as many large-scale datasets focus on exclusively unlicensed or permissively licensed code [5, 3, 2]. The reasons for the exclusion of non-permissively licensed code in other datasets come from potential licensing issues that may be related to the output of models trained on non-permissively licensed data [10]. The Heap is not intended for pre-training models that are aimed at end users, but rather for exclusive use in a research setting. The inclusion of exclusively non-permissively licensed code has the added benefit that it acts as a deterrent for developers to train LLMs on The Heap, ensuring it remains a relevant source of data for downstream tasks. We provide an overview of the licenses used in this work in Table II.
@@ II-C Scraping
For each programming language, we scrape up to 50,000 repositories or as many as are available. Our dataset contains code from repositories created between January 2008 and August 2024. For each selected language, we extract repositories sorted by star count in descending order; this has been used as a loose quality metric before [11]. To maximize extraction efficiency and avoid GitHub’s rate limits, we employ pagination and repository creation date filtering. When the number of repositories within a specified time frame exceeds the rate limit, we narrow the time interval and apply a tumbling window approach to ensure comprehensive coverage. We guide the file extraction based on a list of file extensions from The Stack [5].
@@ III Deduplication
An important aspect of fairly evaluating downstream tasks is preventing data leakage [4]. This is often done through a deduplication process. Although there should be no overlap between our non-permissively licensed dataset and permissively licensed datasets due to our selection procedure, it does not completely prevent overlap [10].
@@
We use a shingle size of $7$ characters, as code files typically use a smaller set of characters compared to large research articles, where $k=9$ [12]. This reduces the likelihood of overly common shingles, which could otherwise inflate similarity scores, as would occur with smaller values of $k$. Files with a Jaccard similarity above $0.7$ are flagged as near duplicates, a threshold shown to be effective for duplicate detection [13].
@@
The Stack V2 [3] All permissively licensed and unlicensed files collected in the Software Heritage [14] archive.
The Stack [1] All permissively licensed repositories collected in the GHArchive [15] and scraped from GitHub.
Red Pajama [2] Repositories from the GitHub dataset hosted by Google BigQuery [16] licensed under MIT, BSD, or Apache licenses.
GitHub Code [8] Repositories from the GitHub dataset hosted by Google BigQuery [16].
CodeParrot [7] All Python files from the GitHub dataset hosted by Google BigQuery [16].
We identify and flag duplicates between our dataset and all publicly available datasets to facilitate a more flexible approach to LLM evaluation, prioritizing both reproducibility and ease of use. This setup minimizes time and computational overhead by removing the burden of duplicate detection from researchers. Users can seamlessly filter data by language or by exact and near-duplicate files, tailoring the dataset to their specific requirements. Table I provides a comprehensive summary of the languages extracted. The third column lists the number of files collected after filtering based on file size and word count. The last column indicates the number of files obtained after removing exact duplicates within our dataset, with exact and near duplicates from other datasets flagged among the remaining files. For more detailed information on the dataset creation process, please refer to the dataset page [^3] https://huggingface.co/datasets/WizzF/Heap-Forge.
@@ III-A Datasets
Our selection of datasets for deduplication is based on previously curated lists [10], with the addition of The Stack V2 [3], which is the only new dataset that has been released since the publication of previous works. We give an overview of all potential datasets in Table III. Due to the comment removal being based on the programming languages of the files, we are not able to infer the correct language for two datasets. The Pile [6], which has been removed and re-uploaded, has lost information about the programming language of a file. Furthermore, due to a known issue with the curation of CodeClippy [^4] https://github.com/CodedotAl/gpt-code-clippy/issues/71, the languages and names of files are misaligned in the dataset. We also exclude this dataset from deduplication. Although we could predict the languages used in the files in these datasets, the tools that provide this functionality do return incorrect predictions, which could result in a duplicate not being removed. As we aim to provide a guarantee that there is no data contamination in our dataset, we remove these two datasets from consideration.
@@ IV Layout
For the file content and metadata, we list the actual content of the file, which is the main information to be used in downstream tasks. We also include information about filename and path, as this has been included in the pre-training procedure of some LLMs [11, 17, 3].
@@ V Future Improvements
We include all files that we scraped that were not duplicates, while this gives us a dataset of deduplicated files, there is still the question of file “quality”. In NLP research, keywords have been used for filtering websites, such as lorem ipsum or TODOs [18], and code datasets have been cleaned of autogenerated files using a similar approach [3]. We believe that this may also affect the quality of code datasets. Specifically, languages that rely heavily on boiler plating, such as Java, may benefit from removing certain common phrases from their corpus. This will be included as a further filtering step in a future release of the dataset.
@@
An under-explored research area involves the presence of multiple natural languages within code. As natural languages are often mixed within one file [19], we plan to adopt a Parts of Speech-like tagging [20] system for the natural languages present in each file. This can give information about the performance of code LLMs when the code is not in English. This will both help the development of non-English code LLMs, as well as aid English-focused LLMs, as they can be evaluated on only English.
@@ VI Limitations/Challenges
In order to use The Heap for a fair evaluation of an LLM, the researcher must be sure that the target LLM has not been trained on The Heap. Aside from our deduplication ensuring this fact for current existing LLMs, our collection process also adds a layer of protection from the inclusion of The Heap in the training procedure. The trend of training LLMs has shifted to only training on permissively licensed data, which would exclude The Heap. Furthermore, the restriction of The Heap to research only, alleviates the problems with author attribution in LLM generations as trained models are not intended to be used by end-users [21, 10].
Furthermore, existing works such as membership inference attacks, have been extended to the scale of entire datasets [22]. This should make it possible in the near future to retroactively test for the inclusion of The Heap in the training procedures of a model.
@@
With the rapid rise of public repositories being used to train code language models, many authors of older repositories were unaware that their code could be utilized for such purposes, leaving them unable to opt-out. Moreover, there is currently no consensus on how developers can opt in or out of having their code included in datasets. We acknowledge these ethical concerns regarding the use of code in deep learning practices and offer the ability for repository owners to opt out of having their code included in our dataset. Although this approach is not ideal, as it places the burden of exclusion on the authors, it aligns with the current best practices [3].
@@ VII Conclusion
We present The Heap, a multilingual dataset of source code that we deduplicated against datasets commonly used in the (pre-)training of large language models. The Heap enables researchers to conduct investigations into the behavior and performance of code large language models without the need to perform extensive deduplication with other datasets. This addresses the shortcomings of LLM investigations not testing for data leakage in 90% of all investigations[4] allowing for more robust conclusions to be made.

### v2

@@ I Introduction
The data-intensive training process of Large Language Models (LLMs) has driven the release of numerous large-scale datasets, particularly for code, to facilitate the development of new models. This rapid increase in the amount of training data used to pre-train LLMs has resulted in extensive datasets covering almost all publicly available code [thestack, together2023redpajama, lozhkov2024starcoder2].
To fairly evaluate LLMs in downstream tasks, fresh data not seen during training is needed. Otherwise such evaluations are contaminated, possibly resulting in overly optimistic results. Unfortunately, obtaining non-contaminated data is increasingly difficult. In fact, a recent study establishes that only 10% of investigations involving LLMs deduplicate their data with respect to the training data in order to avoid contamination [vitale2024catalog].
To address this, we propose The Heap, a dataset of unseen code that can be used for contamination-free evaluation of LLMs in downstream tasks. We address contamination in two ways. First, we select code with a non-permissive license, such as the GNU General Public License. Using such code for training is unattractive, as it may require the end user to publicly release all code in their code bases. Second, we pre-conduct computationally expensive near and exact deduplication, marking code that is used in other datasets widely used for training such as The Stack [thestack].
@@ II Collection
Using the search API, we collect code from GitHub, an online platform for sharing code repositories. This process mimics the collection of other large-scale datasets [kocetkov2023the, lozhkov2024starcoder2, gao2020pile, codeparrot, ghcode], minimizing the probability of including confounding factors in the dataset, such as drifts in the representations of data [bolukbasi2021illusion].
@@ II-A Programming Languages
Ada 676 41,367 34,068
Agda 142 5,483 3,021
ANTLR 101 564 538
Apex 253 17,833 7,561
Assembly 7,100 208,896 101,093
C 50,000 16,585,280 3,076,470
C# 50,000 5,906,716 3,257,456
C++ 50,000 14,891,856 4,469,823
Clojure 27,107 380,567 269,118
Cobol 341 2,242 1,172
Common Lisp 796 45,083 13,922
Coq 477 54,137 22,549
Crystal 368 11,606 6,818
Cuda 1,191 26,948 12,418
D 1,185 185,630 54,034
Dart 11,907 484,935 412,675
EJS 1,475 15,513 12,832
Elixir 2,371 643,856 102,874
Emacs Lisp 377 8,260 7,312
Erlang 1,240 55,932 27,322
F# 876 22,152 13,282
Forth 222 28,287 5,129
Go 50,000 8,506,379 2,328,529
Groovy 2,198 60,299 47,366
Hack 1,379 84,916 37,189
Haskell 8,023 122,788 106,583
Java 50,000 6,989,601 5,168,193
JavaScript 50,000 8,289,901 1,907,803
Julia 2,859 46,284 36,830
Kotlin 21,665 1,467,343 1,042,136
Less 433 17,276 7,308
Lua 42,241 4,605,230 905,120
Mathematica 1,528 164,498 21,208
MATLAB 20,828 1,051,354 599,085
NetLogo 332 900 855
NewLisp 35 5,819 5,123
Nix 1,892 75,093 70,407
Objective-C 7,700 1,899,714 520,332
OCaml 1,961 121,890 60,863
Pascal 5,218 330,832 180,652
Perl 14,673 1,798,520 224,753
PHP 50,000 12,707,727 3,310,243
Processing 2,950 24,723 20,304
Prolog 1,071 38,995 17,570
Python 50,000 2,290,182 1,595,919
R 44,993 589,139 11,679
Raku 158 1,384 689
Ruby 13,378 1,579,655 662,915
Rust 42,847 2,496,177 802,707
Scala 5,893 749,370 210,630
Scheme 1,878 106,620 50,222
Scilab 199 4,531 3,896
SQL 130 47,185 40,800
Starlark 146 524 487
Swift 13,924 633,819 434,849
Vue 14,858 457,605 321,502
WebAssembly 68 834 544
Total 733,663 96,990,250 32,666,778
@@ II-B Query
We focus on repositories that have one of the targeted languages as the main language of the repository. We further select only repositories that are licensed under non-permissive licenses. We choose non-permissive licenses as an initial filter for repositories, as many large-scale datasets focus on exclusively unlicensed or permissively licensed code [kocetkov2023the, lozhkov2024starcoder2, together2023redpajama]. Non-permissively licensed code is usually removed due to potential licensing issues that may be related to the output of models trained on non-permissively licensed data [ourPaper]. The Heap is not intended for pre-training models that are aimed at end users, but rather for exclusive use in a research setting. The inclusion of exclusively non-permissively licensed code has the added benefit that it acts as a deterrent for developers to train LLMs on The Heap, ensuring it remains a relevant source of data for downstream tasks. We provide an overview of the licenses used in this work in Table II.
@@ II-C Scraping
For each programming language, we scrape up to 50,000 repositories or as many as are available. Our dataset contains code from repositories created between January 2008 and August 2024. For each selected language, we extract repositories sorted by star count in descending order; this has been used as a loose quality metric before [li2023starcoder]. To maximize extraction efficiency and avoid GitHub’s rate limits, we employ pagination and repository creation date filtering. When the number of repositories within a specified time frame and star range exceeds the rate limit, we narrow the time interval and apply a tumbling window approach to ensure comprehensive coverage. We guide the file extraction based on a list of file extensions from The Stack [kocetkov2023the].
@@ III Deduplication
An important aspect of fairly evaluating downstream tasks is preventing data leakage [vitale2024catalog]. This is often done through a deduplication process. Although there should be no overlap between our non-permissively licensed dataset and permissively licensed datasets due to our selection procedure, it does not completely prevent overlap [ourPaper].
@@
We use a shingle size of $7$ characters, as code files typically use a smaller set of characters compared to natural language, where $k=9$ [massiveDatasets]. This reduces the likelihood of overly common shingles, which could otherwise inflate similarity scores, as would occur with smaller values of $k$. Files with a Jaccard similarity above $0.7$ are flagged as near duplicates, a threshold shown to be effective for duplicate detection [allamanis].
@@
The Stack V2 [lozhkov2024starcoder2] All permissively licensed and unlicensed files collected in the Software Heritage [softheritage] archive.
The Stack [thestack] All permissively licensed repositories collected in the GHArchive [gharchive] and scraped from GitHub.
Red Pajama [together2023redpajama] Repositories from the GitHub dataset hosted by Google BigQuery [bigquery] licensed under MIT, BSD, or Apache licenses.
GitHub Code [ghcode] Repositories from the GitHub dataset hosted by Google BigQuery [bigquery].
CodeParrot [codeparrot] All Python files from the GitHub dataset hosted by Google BigQuery [bigquery].
We identify and flag duplicates between our dataset and all publicly available datasets to facilitate a more flexible approach to LLM evaluation, prioritizing both reproducibility and ease of use. This setup minimizes time and computational overhead by removing the burden of duplicate detection from researchers. Users can seamlessly filter data by language or by exact and near-duplicate files, tailoring the dataset to their specific requirements. Table I provides a comprehensive summary of the languages extracted. The third column lists the number of files collected after filtering based on file size and word count. The last column indicates the number of files obtained after removing exact duplicates within our dataset, with exact and near duplicates from other datasets flagged among the remaining files. For more detailed information on the dataset creation process, please refer to the dataset page [^3] https://huggingface.co/datasets/AISE-TUDelft/the-heap.
@@ III-A Datasets
Our selection of datasets for deduplication is based on previously curated lists [ourPaper], with the addition of The Stack V2 [lozhkov2024starcoder2], which is the only new dataset that has been released since the publication of previous works. We give an overview of all potential datasets in Table III. Due to the comment removal being based on the programming languages of the files, we are not able to infer the correct language for two datasets. The Pile [gao2020pile], which has been removed and re-uploaded, has lost information about the programming language of a file. Furthermore, due to a known issue with the curation of CodeClippy [^4] https://github.com/CodedotAl/gpt-code-clippy/issues/71, the languages and names of files are misaligned in the dataset. We also exclude this dataset from deduplication. Although we could predict the languages used in the files in these datasets, the tools that provide this functionality do return incorrect predictions, which could result in a duplicate not being removed. As we aim to provide a guarantee that there is no data contamination in our dataset, we remove these two datasets from consideration.
@@ IV Layout
For the file content and metadata, we list the actual content of the file, which is the main information to be used in downstream tasks. We also include information about filename and path, as this has been included in the pre-training procedure of some LLMs [li2023starcoder, team2024codegemma, lozhkov2024starcoder2].
@@ V Future Improvements
We include all files that we scraped that were not duplicates, while this gives us a dataset of deduplicated files, there is still the question of file “quality”. In NLP research, keywords have been used for filtering websites, such as lorem ipsum or TODOs [dodge2021documenting], and code datasets have been cleaned of autogenerated files using a similar approach [lozhkov2024starcoder2]. We believe that this may also affect the quality of code datasets. Specifically, languages that rely heavily on boiler plating, such as Java, may benefit from removing certain common phrases from their corpus. This will be included as a further filtering step in a future release of the dataset.
@@
An under-explored research area involves the presence of multiple natural languages within code. As natural languages are often mixed within one file [pawelka2015code], we plan to adopt a Parts of Speech-like tagging [chiche2022part] system for the natural languages present in each file. This can give information about the performance of code LLMs when the code is not in English. This will both help the development of non-English code LLMs, as well as aid English-focused LLMs, as they can be evaluated on only English.
@@ VI Limitations/Challenges
In order to use The Heap for a fair evaluation of an LLM, the researcher must be sure that the target LLM has not been trained on The Heap. Aside from our deduplication ensuring this fact for current existing LLMs, our collection process also adds a layer of protection from the inclusion of The Heap in the training procedure. The trend of training LLMs has shifted to only training on permissively licensed data, which would exclude The Heap. Furthermore, the restriction of The Heap to research only, alleviates the problems with author attribution in LLM generations as trained models are not intended to be used by end-users [longpre2023data, ourPaper].
Furthermore, existing works such as membership inference attacks, have been extended to the scale of datasets [maini2024llm]. This should make it possible in the near future to test for the inclusion of The Heap in the training procedures of a model.
@@
With the rapid rise of public repositories being used to train code language models, many authors of older repositories were unaware that their code could be utilized for such purposes, leaving them unable to opt-out. Moreover, there is currently no consensus on how developers can opt in or out of having their code included in datasets. We acknowledge these ethical concerns regarding the use of code in deep learning practices and offer the ability for repository owners to opt out of having their code included in our dataset. Although this approach is not ideal, as it places the burden of exclusion on the authors, it aligns with the current best practices [lozhkov2024starcoder2].
@@ VII Conclusion
We present The Heap, a multilingual dataset of source code that we deduplicated against datasets commonly used in the (pre-)training of large language models. The Heap enables researchers to conduct investigations into the behavior and performance of code large language models without the need to perform extensive deduplication with other datasets. This addresses the shortcomings of LLM investigations not testing for data leakage in 90% of all investigations[vitale2024catalog] allowing for more robust conclusions to be made.